OpenStudy (anonymous):

24x^10y/24x^3y^6 simplify

OpenStudy (lgbasallote):

\[\Large \frac{24x^{10}y}{24x^3y^6} \implies \frac{24}{24} \times \frac{x^{10}}{x^3} \times \frac{y}{y^6}\] does that help?

OpenStudy (anonymous):

yes, okay so when I tried doing this problem I got x^7 * y^-5

OpenStudy (lgbasallote):

correct...now simplify it by putting y^5 in the denominator \[\implies \frac{x^7}{y^5}\] got it?

OpenStudy (anonymous):

yes thank you! I am always second guessing myself.

OpenStudy (lgbasallote):

haha welcome ^_^

OpenStudy (anonymous):

ok, so if I rewrite it in an equivalent form with positive exponents would it be x^7/y^5 or do I have to flip flop the equation y^5/x^5

OpenStudy (lgbasallote):

x^7/y^5

OpenStudy (anonymous):

thank you again.

OpenStudy (lgbasallote):

welcome